OpenStudy (anonymous):

A certain radioactive isotope has a half-life of aprox 2000years. How many years would be required for a given amount of this isotope to decay to 60% of that amount?

OpenStudy (anonymous):

\[.6=(\frac{1}{2})^{\frac{t}{2000}}\] solve for t

OpenStudy (anonymous):

\[\frac{t}{2000}=\frac{\ln(.6)}{\ln(.5)}\] \[t=\frac{2000\ln(.6)}{\ln(.5)}\] then a calculator

OpenStudy (anonymous):

Thank u!!